# Vacuum bubble and fissure formation in collective motion with competing attractive and repulsive forces Olivia Clifton1, Angel Chavez2, Antonio Madrigal3, Annie Warren2, Paige Yeung4, and Arnd Scheel2

### v1

@@ 5 Comparison with evolution of finitely many particles
where $\underline{\ell}=[\ell_{1};\ell_{2}]$ is the label for a particle at some position $(x,y)$ in the plane. Modeling repulsion by a Dirac-$\delta$ is no longer meaningful for the ODE and we therefore model the repulsive interaction by a potential $K^{\zeta}(x,y)=(2\pi\zeta)^{-2}K_{0}(\frac{\sqrt{x^{2}+y^{2}}}{\zeta})=\mathcal{F}^{-1}(\frac{1}{4\pi^{2}(1+\zeta^{2}(k^{2}+l^{2}))}),$ for $\zeta$ fixed, small. Here $K_{0}$ represents the modified Bessel function of the second kind. For the attractive part of the potential, we again take $V=2\cos(x)\cos(y)$ for the square lattice, and $V=\cos(x+\frac{y}{\sqrt{3}})+\cos(x-\frac{y}{\sqrt{3}})+\cos(\frac{2y}{\sqrt{3}})$ for the hexagonal lattice.
@@
Onset of Bifurcation: Effect of Repulsive Potential. In addition to finite size corrections, the bifurcation point also changes due to the change in repulsive potential. Replacing the Dirac-$\delta$ by $K^{\zeta}$, and linearizing at the constant state, we find a linearized operator $\mathcal{L}v=\Delta(K^{\zeta}+V)*v.$ The continuum bifurcation points with this potential are thus ${\mu}_{*}=\frac{1}{\pi^{2}(1+2\zeta^{2})}$ (square), and ${\mu}_{*}=\frac{1}{\sqrt{3}\pi^{2}(1+2\zeta^{2})}$ (hexagon), and we do find these in large-$N$ limits in the discrete model below.
@@ 5.1 Numerical Continuation
Fig. 5.2 shows numerically calculated onset of instability of the crystalline state depending on the particle number, compared to the prediction from the continuum limit. We find convergence proportional to the inverse of the total particle number in both square and hexagonal lattice configuration.
@@ 5.2 Direct Simulations and Observed States
Square lattice. On the square lattice, as in the continuum, we find that for a range of $\widetilde{\mu}$} values above onset, perturbations from the crystal state lead to particles forming into fissures. Past a secondary bifurcation point, particles organize only transiently into fissures, which give way into single clusters. In all cases observed, although the crystal state has square particle microstructure, the vacuum states exhibit hexagonal lattice microstructure (with occasional defects as expected in finite-particle simulations). In other words, in the supercritical pattern-forming regime, the predictions from the macrostructure of the problem appear to hold independently of particle microstructure.
@@ 6 Discussion
There are clearly numerous generalizations one might wish to consider. Generalizing to potentials that are not of finite (minimal) rank would be a first step. While the vertical branch still persists and the perturbation argument in §3 should go through, finding solutions with vacuum regions now requires the solution of a free-boundary problem. One may also wish to consider the effect of smoothed repulsive potentials, as used for instance in the discrete setting and analyzed in [13] in the one-dimensional setting.

### v2

@@ 4.4 Vertical Branch II: The Connecting Branch
Remark 4.2.
We emphasize that the bifurcation diagram in the case of the hexagonal lattice, shown in Fig. 4.3, does not include a vertical branch. Bifurcation diagrams with the symmetry of the hexagonal lattice computed in [4] sometimes exhibit such secondary vertical branches, consisting of solutions with $D_{3}$-symmetry. These do not appear to exist in this setting. This is corroborated by numerical calculations of stability, implied in Fig. 4.2 which does not appear to change supercritically.
@@ 5 Comparison with evolution of finitely many particles
where $\underline{\ell}=(\ell_{1},\ell_{2})$ is the label for a particle at some position $(x,y)$ in the plane. Modeling repulsion by a Dirac-$\delta$ is no longer meaningful for the ODE and we therefore model the repulsive interaction by a potential $K^{\zeta}(x,y)=(2\pi\zeta)^{-2}K_{0}(\frac{\sqrt{x^{2}+y^{2}}}{\zeta})=\mathcal{F}^{-1}(\frac{1}{4\pi^{2}(1+\zeta^{2}(k^{2}+l^{2}))}),$ for $\zeta$ fixed, small. Here $K_{0}$ represents the modified Bessel function of the second kind. For the attractive part of the potential, we again take $V=2\cos(x)\cos(y)$ for the square lattice, and $V=\cos(x+\frac{y}{\sqrt{3}})+\cos(x-\frac{y}{\sqrt{3}})+\cos(\frac{2y}{\sqrt{3}})$ for the hexagonal lattice.
@@
Onset of Bifurcation: Effect of Repulsive Potential. In addition to finite size corrections, the bifurcation point also changes due to the change in repulsive potential. Replacing the Dirac-$\delta$ by $K^{\zeta}$, and linearizing at the constant state, we find a linearized operator $\mathcal{L}v=\Delta((K^{\zeta}+V)*v).$ The continuum bifurcation points with this potential are thus ${\mu}_{*}=\frac{1}{\pi^{2}(1+2\zeta^{2})}$ (square), and ${\mu}_{*}=\frac{1}{\sqrt{3}\pi^{2}(1+2\zeta^{2})}$ (hexagon), and we do find these in large-$N$ limits in the discrete model below.
@@ 5.1 Numerical Continuation
Fig. 5.2 shows the numerically calculated onset of instability of the crystalline state depending on the particle number, compared to the prediction from the continuum limit. We find convergence proportional to the inverse of the total particle number in both square and hexagonal lattice configuration.
@@ 5.2 Direct Simulations and Observed States
Square lattice. On the square lattice, as in the continuum, we find that for a range of $\widetilde{\mu}$} values above onset, perturbations from the crystal state lead to particles arranging to form fissures. Past a secondary bifurcation point, particles organize only transiently into fissures, which give way ito single clusters. In all cases observed, although the crystal state has square particle microstructure, the vacuum states exhibit hexagonal lattice microstructure (with occasional defects as expected in finite-particle simulations). In other words, in the supercritical pattern-forming regime, the predictions from the macrostructure of the problem appear to hold independently of particle microstructure.
@@ 6 Discussion
There are clearly numerous generalizations one might wish to consider. Generalizing to potentials that are not of finite (minimal) rank would be a first step, which was carried out for the one-dimensional case in [15]. While the vertical branch still persists and the perturbation argument in §3 should go through, finding solutions with vacuum regions now cannot be reduced to a finite-dimensional problem through an explicit ansatz of the form (2.7), say, and requires parameterization of the free boundary of the support $\partial\Omega_{0}$ in (2.4)–(2.6) by an auxiliary unknown function. One may also wish to consider the effect of smoothed repulsive potentials, as used for instance in the discrete setting and analyzed in [15] in the one-dimensional setting.